OpenStudy (anonymous):

Which of the following has no solution? (x + 1 < -1) ∩ (x + 1 < 1) (x + 1 ≤ 1) ∩ (x + 1 ≥ 1) (x + 1 < 1) ∩ (x + 1 > 1) HEEEELP!

OpenStudy (anonymous):

\[\left( x+1<-1 \right) intersection \left( x+1<1 \right)\] \[\left( x<-2 \right)intersection \left( x<0 \right)\]  similarly check for others.